# Learning to Manipulate under Limited Information

### v2

@@ 3.2.1. Baselines
For baseline comparisons, we consider an agent with full information about a profile and unbounded computational resources, who always picks one of the optimal rankings to submit. We then estimated the average profitability as in (3) of this agent’s submitted ranking across many sampled elections, where as above we continued sampling until the estimated standard error of the mean fell below 5e-4.
3.3. Training Setup
@@
For each voting method $F$, each number $n\in\{5,6,10,11,20,21\}$, each number $m\in\{3,4,5,6\}$, each choice of an input type for the MLP, and each choice of a model size (see the x-axis of Figure 3), we trained one or more “generations” of MLPs with that model size to manipulate elections with $n$ voters and $m$ candidates run using $F$, resulting in over 70,000 trained MLPs. For a given generation, we used the same initialization of MLP weights and the same training, validation, and evaluation profiles for every MLP for $n$ candidates and $m$ voters. Across generations, we varied the initialization of MLP weights and used different training, validation, and evaluation profiles, to provide reassurance that our results were not due to lucky initial weights or profiles. All elections and labels were pre-computed so training could rely fully on the GPU.
@@
We experimented with training the MLPs using different numbers of iterations—between 100 and 1000—different learning rates—1e-3, 3e-3, and 6e-3—and different batch sizes—256 and 512. We hand-tuned these hyper-parameters in order to maximize the performance of all MLPs across all voting methods and profile sizes. For the final training run reported here, we use a batch size of 512 and a learning rate of 6e-3. We train all models for at least 220 iterations and then terminate training with an early stopping rule: after every 20 iterations, we measure the average profitability on a validation batch of 4,096 elections. If 10 validation steps pass without an improvement of at least .001 in average profitability of the submitted ranking, we terminate training. [^9] The average number of training iterations of each MLP was approximately 600 with a standard deviation of approximately 250.
3.3.1. Computing Infrastructure
@@
All code was written in Python using PyTorch and the pref_voting library (pypi.org/project/pref-voting/), version 0.4.42, for all functions dealing with voting. Training and evaluation was parallelized across nine local Apple computers with Apple silicon, the most powerful equipped with an M2 Ultra with 24-core CPU, 76-core GPU, and 128GB of unified memory, running Mac OS 13, as well as up to sixteen cloud instances with Nvidia A6000 or A10 GPUs running Linux Ubuntu 18.04.
4. Results
We begin by discussing our results under the uniform utility model and then turn to the 2D spatial model in Section 4.9.
@@ 4. Results
The average profitability of submitting the optimal ranking in each election (see Baselines above) with different voting methods is shown by the black bars in Figure 1. The other colored bars are for MLP-based manipulators with different types of limited information. Figure 1 shows data for the best performing MLP (for a given number of candidates, number of voters, voting method, and choice of information) with any hidden layer configuration. Figure 3 below shows the performance of MLPs with each of the 26 different hidden layer configuration we tested, focusing on 6 candidates, 10/11 voters, and either the majority matrix or plurality scores choice of limited information. These figures all cover the first generation of trained MLPs. Results for the second and third generations of trained MLPs are qualitatively similar to the first (see Supplementary Material), though we lack a sufficient number of generations to make quantitative statistical claims. The following qualitative highlights are robust across generations. All claims are implicitly qualified to apply to elections with 3–6 candidates and 5-21 voters.
Refer to captionRefer to caption
Figure 1. Results using the uniform probability model. Top: the average profitability of submitted rankings by the best performing MLP with any hidden layer configuration for a given voting method and information type, averaging over 3–6 candidates and 5, 6, 10, 11, 20, and 21 voters. Bottom: the ratio of the average profitability of the MLP’s submitted ranking to that of the ideal manipulator’s submitted ranking.
Refer to captionRefer to caption
@@
Figure 2. Results using the 2D spatial model. Top: the average profitability of submitted rankings by the best performing MLP with any hidden layer configuration for a given voting method and information type, averaging over 3–6 candidates and 5, 6, 10, 11, 20, and 21 voters. Bottom: the ratio of the average profitability of the MLP’s submitted ranking to that of the ideal manipulator’s submitted ranking.
4.1. The importance of majorities
Sufficiently large MLPs learned to profitably manipulate all voting methods we studied on the basis of knowing only the majority matrix, though the profitability of such manipulation varied dramatically across methods. Interestingly, we did not find a substantial increase in profitability of manipulation for MLPs that learned to manipulate on the basis of the more informative margin matrix instead of the majority matrix, except in the case of Borda (especially for 6 candidates, as shown in the Supplementary Material). In fact, the qualitative margin matrix was about as useful as the margin matrix for learning to manipulate Borda.
4.2. The limited usefulness of plurality scores
Though obviously useful for manipulating Plurality and somewhat useful for several methods for 3 candidates (though less useful for 4 or 5 candidates), knowing the plurality scores was insufficient in 6-candidate elections for profitably manipulating methods other than Plurality (though Borda may be barely manipulable in this case), as shown in Figure 3. Moreover, in the case of manipulating Plurality, learning to manipulate on the basis of the plurality ranking (see Figure 1) led to profitability comparable to learning to manipulate on the basis of the plurality scores themselves.
4.3. Highly manipulable vs. resistant methods
Plurality and especially Borda have long been regarded as highly manipulable. [^10] Jean-Charles de Borda declared that his method was “intended for only honest men” Black (1958). On the basis of heuristic algorithms for manipulating Borda, Walsh (2011) concludes that “Borda voting can usually be manipulated with relative ease” (p. 13). For empirical results on manipulation of Borda by humans, see Kube and Puppe (2009). Our results show that this is so even under limited information, e.g., in the form of the majority matrix.
IRV and IRV-PUT were quite resistant to manipulation on the basis of limited information (with the exception of the manipulability of IRV for 3 candidates and 10 voters), despite the fact that these methods are more manipulable than some others by an ideal manipulator. In addition, Minimax and Split Cycle stood out for their resistance to manipulation. It it is noteworthy that while Minimax and Split Cycle were not much more profitably manipulable than IRV and IRV-PUT on the basis of the majority matrix, qualitative margin matrix, or margin matrix in absolute terms (see top of Figure 1), MLPs came closer to the ideal manipulator for learning to manipulate Minimax and Split Cycle on the basis of this information than for IRV and IRV-PUT (see bottom of Figure 1), which are more manipulable by an ideal manipulator. Another noteworthy difference is that IRV and IRV-PUT are fairly manipulable by a voter with only the sincere winners information (see Figure 1), whereas Minimax and Split Cycle are not.
4.4. The subtleties of tiebreaking
While the distinction between IRV and IRV-PUT is subtle, it leads to significant differences with respect to single-voter manipulability under 3 candidates and an even number of voters (see graphs in the Supplementary Material). Of course, manipulation by a single voter is only possible in very close elections, in which case ties matter.
4.5. Parity of the number of voters
Indeed, the parity of the number of voters is a key factor for some methods. This is most striking for Stable Voting, which is barely manipulable with 11 voters but more manipulable with 10 (see Figure 3). A key difference is that with 10 voters, it is possible to have margins of zero between candidates, in which case a single voter has more manipulative power under Stable Voting, which produces fewer tied elections in the presence of zero margins than other methods like Minimax.
4.6. Effect of the number of candidates
For an ideal manipulator, manipulation becomes more profitable as the number of candidates increases (see the Supplementary Material). The same can be true for the MLP-based manipulators up to a point, e.g., 4 or 5 candidates for most voting methods when using the majority, qualitative margin, or margin information. However, the ratio between the profitability of the MLP-based manipulator’s submitted rankings and those of the ideal manipulator declines as the number of candidates increases from 3 to 6 (see Supplementary Material). This is intuitive, as manipulation should increase in complexity with more candidates.
4.7. Profitability and ease of learnability
Not only is it more profitable to manipulate, say, Borda than Stable Voting, but it is also easier to learn—in the sense of requiring smaller MLPs: whether we consider one, two, or three hidden layers, in each case a smaller network is needed to learn to profitably manipulate Borda on the basis of the majority matrix (see Figure 3).
4.8. Worst-case complexity vs. learnability
Despite the NP-hardness of deciding if one can manipulate Nanson so as to elect a desired candidate, it is still possible to learn to do so on the basis of just the majority matrix for elections with up to 6 candidates. In this connection, studying manipulability with more candidates is of great interest.
4.9. Differences with the 2D spatial model
Comparing Figure 1 for the uniform utility model and Figure 2 for the 2D spatial model model, the most striking differences are (1) that all voting methods become less profitably manipulable (roughly by one half) under the spatial model and (2) even the best MLP’s could not learn to profitably manipulate against Minimax, Nanson, and Split Cycle under the spatial model. [^11] A natural thought to explain (2) is that Minimax, Nanson, and Split Cycle are Condorcet consistent, and there is a high frequency of Condorcet winners under the 2D spatial model (yet this must be squared with the results for Stable Voting). However, even when there is a Condorcet winner in $\mathbf{P}$ and we are using a Condorcet voting method, a voter may still have an incentive to submit an insincere ranking in order to create a majority cycle, possibly resulting in a different winner. On the other hand, such possibilities are evidently rare and difficult to learn to exploit. On the other hand, the comparative usefulness, for manipulating against each voting method, of the different types of limited information is largely the same under the spatial model as under the uniform utility model (this is true even for Minimax, Nanson, and Split Cycle, looking at which types of information produce less negative results). We conjuncture that these findings about types of limited information are robust across other standard probability models as well.
@@ 4.9. Differences with the 2D spatial model
Figure 3. Results using the uniform probability model with 6 candidates and 10/11 voters for MLPs manipulating on the basis of the plurality scores or the majority matrix. Error bars indicate twice the estimated standard error of the mean. Hidden layer configurations of trained MLPs are shown on the x-axis.
5. Conclusion
@@
In committee-sized elections (5-21 voters), MLPs can learn to vote strategically on the basis of limited information, though the profitability of doing so varies significantly between different voting methods. This serves as a proof of concept for the study of machine learnability of manipulation under limited information. There are a number of natural extensions for future work, including manipulation by a coalition of multiple voters, as well as different probability models for generating elections. Our code is already set up to handle these extensions, which only require more compute. However, further research is needed on other questions: What if all agents in the election strategize? And what is the social cost or benefit of the learned manipulations? Finally, one limitation of the classification approach in this paper is that it is infeasible to apply to more than 6 candidates. To overcome this limitation, we plan to develop a reinforcement learning approach to learning to manipulate.
Acnowledgements
@@
For helpful feedback, we thank Dominic Hughes, the anonymous reviewers for SCaLA, and the participants in the Values-Centered AI seminar at the University of Maryland in Fall 2023, the logic seminar at the University of Maryland in Spring 2024, and the Interdisciplinary Workshop on Computational Social Choice at the Center for Human-Compatible Artificial Intelligence in Spring 2024.

### v3

@@ Introduction
For baseline comparisons, we consider an agent with full information about a profile and unbounded computational resources, who always picks one of the optimal rankings to submit. We estimated the average profitability as in (4) of this agent’s submitted ranking across many sampled elections, where as above we continued sampling until the estimated standard error of the mean fell below 5e-4.
Training Setup
@@
For each voting method $F$, each $n\in\{5,6,10,11,20,21\}$, each $m\in\{3,4,5,6\}$, each choice of an input type for the MLP, and each choice of a model size (see the $x$-axis of Figure 2), we trained one or more “generations” of MLPs with that model size to manipulate elections with $n$ voters and $m$ candidates run using $F$, resulting in over 100,000 trained MLPs. For a given generation, we used the same initialization of MLP weights and the same training, validation, and evaluation profiles for every MLP for $n$ candidates and $m$ voters. Across generations, we varied the initialization of MLP weights and used different training, validation, and evaluation profiles, to provide reassurance that our results were not due to lucky initial weights or profiles. All elections and labels were pre-computed so training could rely fully on the GPU.
@@
We experimented with training the MLPs using different numbers of iterations—between 100 and 1000—different learning rates—1e-3, 3e-3, and 6e-3—and different batch sizes—256 and 512. We hand-tuned these hyper-parameters to try to maximize the performance of all MLPs across all voting methods and profile sizes. For the final training run reported here, we use a batch size of 512 and a learning rate of 6e-3. We train all models for at least 220 iterations and then terminate training with an early stopping rule: after every 20 iterations, we measure the average profitability on a validation batch of 4,096 elections. If 10 validation steps pass without an improvement of at least .001 in average profitability of the submitted ranking, we terminate training. [^10] The average number of training iterations of each MLP was approximately 600 with a standard deviation of approximately 250.
Computing Infrastructure
@@
All code was written in Python using PyTorch, version 2.0.1, and the pref_voting library (pypi.org/project/pref-voting/), version 0.4.42 or later. Training and evaluation was parallelized across nine local Apple computers with Apple silicon, the most powerful equipped with an M2 Ultra with 24-core CPU, 76-core GPU, and 128GB of unified memory, running Mac OS 13, as well as up to sixteen cloud instances with Nvidia A6000 or A10 GPUs running Linux Ubuntu 18.04.
Results
@@
The average profitability of submitting the optimal ranking in each election (see Baselines above) with different voting methods is shown by the black bars in Figure 1. The other colored bars are for MLP-based manipulators with different types of limited information. Figure 1 shows data for the performance (averaging over different numbers of candidates and voters) of the best performing MLPs (for each number of candidates, number of voters, voting method, and choice of information) with any hidden layer configuration.
Figure 2 and Supplementary Figures A.2-B.3 show the performance of MLPs with each of the 26 different hidden layer configurations we tested, focusing on 6 candidates, 10/11 voters, and different choices of limited information. All Supplementary Figures cited are available at github.com/epacuit/ltm/blob/main/supplementary-figures.pdf.
These figures all cover the first generation of trained MLPs. Results for the second and third generations of trained MLPs are qualitatively similar to the first (see the GitHub repository), though we lack a sufficient number of generations to make quantitative statistical claims. [^11] However, we can make claims of statistical significance about the differences in performance between two particular, trained MLPs, e.g., one trained using the majority matrix vs. one trained using the plurality scores (see the GitHub repository). The following qualitative highlights are robust across generations. All claims are implicitly qualified to apply to elections with 3–6 candidates and 5-21 voters.
Uniform Utility Model
Mallows Model
Spatial 2D Model
@@
Figure 1: Left: the average profitability of submitted rankings by the best performing MLP with any hidden layer configuration for a given voting method and information type, averaging over 3–6 candidates and 5, 6, 10, 11, 20, and 21 voters. Right: the ratio of the average profitability of the MLP’s submitted ranking to that of the ideal manipulator’s submitted ranking.
@@
Figure 2: Results using the uniform utility model with 6 candidates and 10/11 voters for MLPs manipulating on the basis of the plurality scores or majority matrix. Error bars indicate twice the estimated standard error of the mean. Hidden layer configurations of trained MLPs are shown on the x-axis. Versions of this figure for the Mallows model, spatial 2D model, and different types of limited information appear in Supplementary Figures A.2-B.3.
Differences across probability models
Comparing the results for the three probability models in Figure 1, we see that MLPs achieved the highest profitability in manipulating against voting methods in elections generated by the uniform utility model, followed by the Mallows model, followed by the spatial 2D model. While the graphs for Mallows are roughly scaled down versions of the those for the uniform model, it is striking that for the spatial 2D model, even the best MLP’s could not learn to profitably manipulate against Minimax, Nanson, and Split Cycle. [^12] A natural thought to explain this is that Minimax, Nanson, and Split Cycle are Condorcet consistent, and there is a high frequency of Condorcet winners under the spatial 2D model (yet this must be squared with the results for Stable Voting). However, even when there is a Condorcet winner in $\mathbf{P}$ and we are using a Condorcet voting method, a voter may still have an incentive to submit an insincere ranking in order to create a majority cycle, possibly resulting in a different winner. On the other hand, such possibilities are evidently rare and difficult to learn to exploit. On the other hand, the comparative usefulness, for manipulating against each voting method, of the different types of limited information is largely the same under all three models (this is true even for Minimax, Nanson, and Split Cycle under the spatial model, looking at which types of information produce less negative results). We conjecture that these findings about types of limited information are robust across other standard probability models as well.
The importance of majorities
With the uniform utility and Mallows models, sufficiently large MLPs learned to profitably manipulate all eight voting methods on the basis of knowing only the majority matrix, though the profitability of such manipulation varied dramatically across methods. Interestingly, we did not find a substantial increase in profitability of manipulation for MLPs that learned to manipulate based on the more informative margin matrix instead of the majority matrix, except in the case of Borda and Black’s (especially for 6 candidates, as shown in Supplementary Figures B.1-B.3). In fact, the qualitative margin matrix was about as useful as the margin matrix for learning to manipulate Borda and Black’s.
The limited usefulness of plurality scores
Though obviously useful for manipulating Plurality and somewhat useful for several methods for 3 candidates (though less useful for 4 or 5), knowing the plurality scores was insufficient in 6-candidate elections for profitably manipulating methods other than Plurality (though Borda may be barely manipulable in this case), as shown in Figure 2 and Supplementary Figures A.2-A.3. Moreover, in the case of manipulating Plurality, learning to manipulate on the basis of the plurality ranking led to profitability comparable to learning on the basis of the plurality scores themselves (see Figure 1).
Highly manipulable vs. resistant methods
Plurality and especially Borda have long been regarded as highly manipulable. [^13] Borda declared that his method was “intended for only honest men” (Black 1958). Based on heuristic algorithms for manipulating Borda, Walsh (2011) concludes that “Borda voting can usually be manipulated with relative ease” (p. 13). For empirical results on manipulation of Borda by humans, see Kube and Puppe 2009. Our results show that this is so even under limited information, e.g., the majority matrix. The manipulability of Borda seems to infect Black’s method as well, as it uses Borda when there is no Condorcet winner.
IRV-PUT was quite resistant to manipulation on the basis of limited information, despite the fact that it is more manipulable than some others by an ideal manipulator. In addition, Minimax and Split Cycle stood out for their resistance to manipulation, especially under the spatial 2D model. For the uniform utility and Mallows models, it is noteworthy that while Minimax and Split Cycle were not much more profitably manipulable than IRV-PUT on the basis of the majority matrix, qualitative margin matrix, or margin matrix in absolute terms, MLPs came closer to the ideal manipulator for learning to manipulate Minimax and Split Cycle based on this information than for IRV-PUT (see Figure 1), which is more manipulable by an ideal manipulator. Another noteworthy difference is that an MLP with only the sincere winners information can achieve between 30-40% of the profitability of an ideal manipulator when manipulating against IRV-PUT (see Figure 1), whereas the average profitability of manipulation by an MLP with sincere winners information is negative or barely positive when manipulating against Minimax or Split Cycle.
The subtleties of tiebreaking
As noted in Footnote 3, there are different ways of dealing with ties in first-place votes for Instant Runoff. This actually leads to significant differences with respect to single-voter manipulability under 3 candidates and an even number of voters (see Supplementary Figure C.1). Of course, manipulation by a single voter is only possible in very close elections, in which case ties matter.
Parity of the number of voters
The parity of the number of voters is a key factor for some methods. This is most striking for Stable Voting (also see Black’s), which is barely manipulable with 11 voters but more manipulable with 10 (see Figure 2 and Supplementary Figures A.2-A.3). A key difference is that with 10 voters, it is possible to have margins of zero between candidates, in which case a single voter has more manipulative power under Stable Voting, which produces fewer tied elections in the presence of zero margins than other methods like Minimax.
Figure 3: Top: average profitability of submitted rankings by an ideal manipulator. Middle: average profitability by the best performing MLP with any hidden layer configuration using the majority matrix information, averaging over 5, 6, 10, 11, 20, and 21 voters. Bottom: the ratio of the average profitability of the MLP’s submitted ranking to the average profitability of the ideal manipulator’s submitted ranking. Versions of this figure for the Mallows model and the spatial 2D model appear in Supplementary Figures D.2-3.
Effect of the number of candidates
For an ideal manipulator, manipulation becomes more profitable as the number of candidates increases (see Figure 3 and Supplementary Figures D.2-D.3). The same can be true for the MLP-based manipulators up to a point, e.g., 4 or 5 candidates for most voting methods when using the majority, qualitative margin, or margin information. However, the ratio between the profitability of the MLP-based manipulator’s submitted rankings and those of the ideal manipulator declines as the number of candidates increases from 3 to 6 (again see the figures just cited). This is intuitive, as manipulation should increase in complexity with more candidates.
Profitability and ease of learnability
Not only is it more profitable to manipulate, say, Borda than Stable Voting, but it is also easier to learn—in the sense of requiring smaller MLPs: whether we consider one, two, or three hidden layers, a smaller network is needed to learn to profitably manipulate Borda compared to Stable Voting on the basis of the majority or margin matrix (see Figure 2 and Supplementary Figures A.2-B.3).
Worst-case complexity vs. learnability
Despite the NP-hardness (when we allow the number of candidates to increase) of deciding if one can manipulate Nanson so as to elect a desired candidate, it is still possible (under the uniform utility and Mallows models) to learn to manipulate Nanson to achieve an increase in expected utility. In this connection, it would be interesting to study learning to manipulate with more candidates.
Conclusion
@@
In committee-sized elections (5-21 voters), MLPs can learn to vote strategically on the basis of limited information, though the profitability of doing so varies significantly between different voting methods. This serves as a proof of concept for the study of machine learnability of manipulation under limited information. There are a number of natural extensions for future work, including manipulation by a coalition of multiple voters, additional probability models for generating elections, and training an MLP to manipulate in elections with varying numbers of voters and candidates (rather than training different MLPs for each choice of a number of voters and number of candidates). Our code is already set up to handle these extensions, which only require more compute. However, further research is needed on other questions: What if all agents in the election strategize? What is the social cost or benefit of the learned manipulations? How do different neural network architectures affect results? Finally, one limitation of the classification approach in this paper is that it is infeasible to apply to more than 6 candidates. To overcome this limitation, we plan to develop a reinforcement learning approach to learning to manipulate.
Acknowledgments
@@
For helpful feedback, we thank Dominic Hughes, the anonymous reviewers for SCaLA 2024 and for AAAI 2025, and the participants in the Values-Centered AI seminar at the University of Maryland in Fall 2023, the logic seminar at the University of Maryland in Spring 2024, the Interdisciplinary Workshop on Computational Social Choice at the Center for Human-Compatible Artificial Intelligence in Spring 2024, and New Directions in Social Choice at EC 2024.